OpenStudy (anonymous):

Take the derivative of y=(1-cosx)/(sin^2x)

OpenStudy (anonymous):

So far I got this, and I'm not sure if the first is correct or not \[y=\frac{ -(-sinx)(\sin^2x)-(1-cosx)2sinx(cosx)}{ (sinx)^4 }\] ???

OpenStudy (anonymous):

I used the quotient rule and the chain rule

OpenStudy (freckles):

That is looks good

OpenStudy (freckles):

I think I can make your problem simpler though \[y=\frac{1-\cos(x)}{1-\cos^2(x)}=\frac{1-\cos(x)}{(1-\cos(x))(1+\cos(x))}=\frac{1}{1+\cos(x)}=(1+\cos(x))^{-1}\] if you think that is simpler

OpenStudy (freckles):

and then differentiate but what you have is good so far

Nnesha (nnesha):

http://prntscr.com/73jtkp sorry to interrupt hmm

OpenStudy (anonymous):

My teacher wants us to use the original question... I got stuck in simplfying it more. I tried common factoring the sinx out and then I got stuck...

OpenStudy (freckles):

\[y'=\frac{\sin^3(x)-(1-\cos(x))2\sin(x)\cos(x)}{\sin(x)(\sin(x))^3} \\ y'=\frac{\sin(x)[\sin^2(x)-(1-\cos(x))2\cos(x)]}{\sin(x)(\sin(x))^3}\] so if I write this does this lose you?

OpenStudy (anonymous):

Okay, I got that

OpenStudy (freckles):

then we cancel out the common factor sin(x) on top and bottom (and yes trig everywhere tonight @Nnesha )

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

\[y'=\frac{\sin^2(x)+(\cos(x)-1)2 \cos(x)}{(\sin(x))^3} \\ y'=\frac{\sin^2(x)+2\cos^2(x)-2\cos(x)}{\sin^3(x)} \\ y'=\frac{\sin^2(x)+\cos^2(x)+\cos^2(x)-2\cos(x)}{\sin^3(x)}\] can you think of any identity you can apply to the numerator right now?

OpenStudy (anonymous):

sin^2x+cos^2x=1 ?

OpenStudy (freckles):

yea

OpenStudy (freckles):

\[y'=\frac{\cos^2(x)-2\cos(x)+1}{\sin^3(x)}\] you can factor the numerator

OpenStudy (anonymous):

Oh I got it now! Thank you so much! :D I forgot to use my identities.

OpenStudy (freckles):

oh is that the way their final answer looks ?

OpenStudy (anonymous):

It's more simplify: \[y'=\frac{ (cosx-1)^2 }{ \sin^3x }\]

OpenStudy (freckles):

ok cool we can actually "simplify it further " \[y'=\frac{(\cos(x)-1)^2}{\sin^2(x) \sin(x)} =\frac{(\cos(x)-1)^2}{(1-\cos^2(x))\sin(x)} \\ y'=\frac{(\cos(x)-1)(\cos(x)-1)}{(1-\cos(x))(1+\cos(x))\sin(x)} \\ y'=\frac{-(\cos(x)-1)}{(1+\cos(x) )\sin(x)} \\ y'=\frac{1-\cos(x)}{(1+\cos(x))\sin(x)} \cdot \frac{1+\cos(x)}{1+\cos(x)} \\ y'=\frac{1-\cos^2(x)}{(1+\cos(x))^2(\sin(x)}=\frac{\sin^2(x)}{(1+\cos(x))^2\sin(x)} \\ y'=\frac{\sin(x)}{(1+\cos(x))^2}\] depending on what "more simplified" means to you :p

OpenStudy (freckles):

the only thing you have have to be careful about is the domain thing

OpenStudy (freckles):

my y' shouldn't exist at x=npi where n is an integer

OpenStudy (anonymous):

Wow! I didn't realized that you could actually simplified even more! Thank you for showing me that! :)

OpenStudy (freckles):

no problem that is also what you get when you differentiate the simplified form of your function (though when you simplify functions you may be messing with the domain of the original function\[y=(1+\cos(x))^{-1} \\ y=(-1)(1+\cos(x))^{-2}(-\sin(x))=\frac{\sin(x)}{(1+\cos(x))^2}\])

OpenStudy (cookie_2046):

:)

OpenStudy (freckles):

anyways peace and have fun